# ScNeuGM: Scalable Neural Graph Modeling for Coloring-Based Contention and Interference Management in Wi-Fi 7

### v1

@@ VI-E1 Comparison with PG and DPG
We compare the performance of the ES with the PG and the DPG when a fixed small number of STAs is selected using the batching mode of the DHF, i.e., $K^{\prime}=20$. The performance is measured based on the QoS violations, the ratio of approximated optimal slots to the number of used slots, and the averaged reward performance indicator. The results show that the proposed ES algorithm achieves the lowest QoS violation probability, the least number of slots, and the highest reward performance (approximately $4\sim 10$ times higher), while the PG and DPG algorithms struggle to converge to satisfactory performance. This is because the PG and DPG algorithms require STA-pairwise (or edge-wise) gradients, but no STA-pairwise feedback is available. In comparison, the ES algorithm directly applies the reward feedback in the parameter space to improve the NGM NN.
Figure 11: The performance of ES, PG and DPG for the NGM when $K^{\prime}=20$.
@@ VI-E2 Performance of ES with Different Batching Schemes
We then show the performance of the ES algorithm with and without the batching mode of the DHF. Specifically, we consider 1) the proposed batching method that adaptively increases the number of STAs, $K^{\prime}$, based on the reward performance (with legend “Adp-B”), as shown in Algorithm 2; 2) a batching method that linearly increases $K^{\prime}$ by one in each step (with the legend “Lin-B”); and 3) a method that uses no batching but trains the NGM NN with the total number of STAs, $K=K^{\prime}$ (with the legend “No-B”). Fig. 12 measures the number of STAs, the training time, and the averaged reward performance indicator during the training steps. The results show that the proposed adaptive batching converges to the target number of STAs at approximately $1100$ steps. Additionally, the performance of all three compared methods is similar around $1100$ steps. However, the training time of the proposed adaptive batching method is approximately $4$ times and $2$ times shorter than that of the linear batching method and the no-batching method, respectively. In other words, the proposed method significantly reduces the training time for the NGM. In the remaining simulations, the trained NGM parameters are fixed as those from the last step of the proposed ES with the adaptive batching, without further updates.
VI-F Performance of NGM with Bucketing Mode of DHF
We evaluate the trained NGM NN online with the bucketing mode of DHF when there are $K=1000$ STAs.
Figure 13: The NGM NN in the architecture when STAs are static.
Figure 14: Comparison between the NGM and heuristic graph models.
@@ VI-F1 Performance When STAs are Static
Fig. 13 shows the NGM NN’s performance in the cumulative distribution function (CDF) of the number of used slots and the number of QoS violated STAs when the NGM NN operates for different numbers of rounds, $m$, in the architecture. In this architecture, all STA pairs connected by edges in previous rounds are combined into the processing STA pairs for the current round, as shown in (35). The results show that across different rounds, the number of slots to color the generated graph remains consistent, around $30$ slots. It also shows that the NGM NN in the first round results in $2\%$ of STAs experiencing QoS violations. However, by running the NGM process for more rounds, the QoS violations are fewer than $1\%$. This improvement is due to the bucketing mode in each round, which can miss some contending and interfering STA pairs, as previous simulations in Section VI-D. By adding the previously connected pairs in the NGM process as (35), the likelihood of missing the processing of contending or interfering STA pairs is efficiently reduced.
Fig. 14 compares the number of slots and QoS violations achieved by the NGM with the DHF’s bucketing mode when the number of rounds $m$ is 9 (with legend “NGM-B”), the NGM with all STA pairs processed (with legend “NGM-A”), and the heuristic graph models (with legends “CHG” and “IFG”). The results show that the NGM achieves $25\%$ fewer slots than the heuristic graph models. This is because the NGM allows slot sharing among STA pairs that are not heavily contending or interfering with each other, while the heuristic graph models separate all contending or interfering pairs. Consequently, the heuristic graph models achieve fewer QoS violations but use more slots or cause higher delays in periodicity. The results also show that the scheme with partly processed STA pairs using the DHF’s bucketing mode achieves performance similar to the one where the NGM processes all STA pairs. This implies that only a subset of STA pairs, e.g., those likely to be contending or interfering, requires NGM processing, which further suggests that the DHF’s bucketing mode reduces the complexity when modeling the graph.
@@ VI-F2 Performance When STAs are Mobile
We then show the NGM’s performance when STAs are mobile. Specifically, STAs move in a random direction within the simulated area, with speeds varying from 00 to $5$ meters per second, approximately the typical walking speeds of mobile robots. Note that when STAs reach the boundary of the simulated area, they move in another random direction toward the inside of the area. We compare the following schemes: 1) the NGM using the DHF’s bucketing mode and combining previous rounds’ edges (the maximum rounds to combine is $I=20$) into the processed pairs (with legend “NGM-B”); 2) the NGM using the DHF’s bucketing mode without combining previous edges in the NGM processed pairs (with legend “No-Comb”); and 3) the NGM processing all edge pairs among STAs (with legend “NGM-A”). Fig. 15 measures the average number of slots and the average packet loss rates in the above schemes. The results show that the average number of slots is similar across all schemes (around $30$ slots), while the NGM-B and NGM-A schemes achieve lower packet loss rates. This is because the NGM-B and NGM-A schemes process more contending or hidden STA pairs than the No-Comb scheme. On the other hand, the NGM-B scheme has $30\%$ fewer packet losses than the NGM-A scheme when STAs have high mobility. This is because the DHF’s bucketing mode selectively processes contending or interfering STA pairs, making the NGM process with the DHF’s bucketing mode faster when modeling the graph. Consequently, it responds more timely to the varying locations of STAs. The detailed computing time of the NGM process is measured next.
@@ VI-G Computational Complexity in Time
We discuss the computational complexity of the proposed methods in terms of the number of STAs, $K$. Training the state embedding (trained over STA-wise), the predictors (trained over STA-pairwise), and the DHF (also trained over STA-pairwise) incurs $\mathcal{O}(K)$, $\mathcal{O}(K^{2})$, and $\mathcal{O}(K^{2})$ complexity, respectively. The batching mode of the DHF requires a $\mathcal{O}(K)$ complexity to compare the random query and collect the batch. Training the EGNN in Algorithm 2 is also processed over STA-pairwise, resulting in an $\mathcal{O}(K^{2})$ complexity. The NGM complexity when modeling/processing the graph depends on the number of STA pairs collected using the bucketing mode of the DHF. Since it is difficult to analyze hash codes/tables generated by DHF, we instead measure the computing time of the NGM process quantitatively. Specifically, Fig. 16 compares the computing time of the NGM process when $K=1000$ in 1) the scheme where the NGM only processes STA pairs collected by the DHF bucketing mode (with legend “NGM-B”), and 2) the scheme where the NGM processes all STA pairs (with legend “NGM-A”). The results show that the DHF’s bucketing mode reduces the computing time for STA-pairwise operations, such as predictors and edge generation, by a factor of $8$, and the total computing time by a factor of $3$. This is because only the STA pairs in the buckets are processed, while the bucketing mode adds some additional computing time when collecting the STA pairs in the buckets/hash tables. Note that the current hash table construction is single-threaded on a central processing unit, while further time reduction can be achieved by an efficient parallelized implementation on a graphics processing unit.
Figure 15: The NGM performance in online architecture when STAs are mobile.
Figure 16: The computing time, $\operatorname{\mathbb{E}}[\tau^{(m)}]$, for the NGM process to model and color the graph, including the time spent by the state embedding (Emb.), the DHF’s hash code generation (Hsh.), the DHF’s bucketing mode (Buc.), the predictors (Pre.), edge/graph generation (EG) and the graph coloring (Col.)
@@ VII Conclusion
This paper presented ScNeuGM for managing contention and interference in Wi-Fi 7 networks through coloring-based slot assignment. ScNeuGM utilized the ES algorithm, which eliminates the reliance on explicit STA-pairwise feedback in the NGM, offering a robust and scalable optimization approach. Furthermore, the ScNeuGM leveraged a DHF to group contending or hidden STAs and reduce the computational complexity of training and inference of the NGM NN. Extensive evaluation demonstrated the scalability and effectiveness of our methods in large Wi-Fi network scenarios. Future work could focus on further optimizing the graph coloring algorithm and accelerating the computation routines in the NGM. Additionally, research could be conducted to explore richer graph representations that integrate the application layer’s context, semantic, and task-oriented information, and to extend the framework to accommodate more heterogeneous network deployments.
Appendix: Proof of Proposition 1
Proof.
To prove this statement, we construct an example that satisfies it. Let $\hat{\mathbf{z}}$ and $\hat{Z}$ be an optimal solution of problem (9). In the adjacency matrix $\hat{\mathbf{E}}$, we set $\hat{E}_{i,j}=1$ for every STA pair $(i,j)$ with different colors and set $\hat{E}_{i,j}=0$ otherwise. In other words, $\hat{\mathbf{E}}$ represents a complete multipartite graph. The graph with $\hat{\mathbf{E}}$ has a unique coloring scheme $\hat{\mathbf{z}}$ and $\hat{Z}$ because any alternative coloring of STAs would cause violations of the edge constraints. Therefore, $\hat{\mathbf{z}},\hat{Z}=\chi(\hat{\mathbf{E}})$, which also solves the contention and interference management problem according to the definition of $\hat{\mathbf{z}}$ and $\hat{Z}$. By providing this example, we demonstrate that an optimal graph exists, thereby proving the statement. ∎
Appendix: Proof of Theorem 1
Proof.
Given the optimal graph $\mathcal{G}^{*}$ with the adjacency matrix $\mathbf{E}^{*}$ and its coloring scheme $\mathbf{z}^{*},Z^{*}=\chi(\mathbf{E}^{*})$, the coloring scheme of the optimal graph, $\mathbf{z}^{*}$ and $Z^{*}$, corresponds to the optimal slot assignment, where all STAs’ reliability constraints are satisfied, and the number of slots is minimized. Suppose we remove all edges between non-contending and non-hidden STAs in $\mathcal{G}^{*}$. This results in a new graph $\mathcal{G}^{\prime}$ with edge values $E^{\prime}_{i,j}=E^{*}_{i,j}\land(O^{\mathrm{C}}_{i,j}+O^{\mathrm{H}}_{i,j})$ for all $i\neq j$. By re-coloring $\mathcal{G}^{\prime}$, STAs in $\mathcal{G}^{\prime}$ will share the same color (or the same slot) if and only if they have no edges between them in $\mathcal{G}^{\prime}$, i.e., they are non-contending and non-hidden STAs, or they were in the same color in $\mathcal{G}^{*}$. In this new color scheme, i.e., $\mathbf{z}^{\prime},Z^{\prime}=\chi(\mathbf{E}^{\prime})$, each STA now shares the slot only with: 1) STAs that are neither contending with nor hidden from it, or 2) those STAs that were sharing slots in the previous optimal graph $\mathcal{G}^{*}$. Note that each STA’s transmissions are not affected by non-contending and non-hidden STAs, and the slot-sharing STAs in the previous graph $\mathcal{G}^{*}$ have reliable transmissions, according to the definition of its optimality. Therefore, the coloring scheme of the new graph $\mathcal{G}^{\prime}$, i.e., $\mathbf{z}^{\prime}$ and $Z^{\prime}$, ensures reliable transmissions, thus satisfying the constraints. Note that $\mathcal{G}^{\prime}$ is generated by edge removal from $\mathcal{G}^{*}$. Thus, their chromatic numbers satisfy $Z^{\prime}\leq Z^{*}$. Further, according to the optimality of $\mathcal{G}^{*}$, $Z^{*}$ is the minimum number of slots required to ensure reliability, implying that $Z^{\prime}$ is equal to $Z^{*}$. Therefore, the coloring scheme of $\mathcal{G}^{\prime}$, i.e., $\mathbf{z}^{\prime}$ and $Z^{\prime}$, ensures the reliability constraints and minimizes the number of slots. In other words, $\mathcal{G}^{\prime}$ is optimal, which proves the statement.
Appendix: Pre-training of NGM NN
The SENN is trained as an encoder NN that maps the STA state sequence into an embedding vector, along with a decoder NN to recover the sequence from the embedding vector. The decoder $\mathrm{DCNN}(\cdot|\theta^{\mu}_{\mathrm{DC}})$ has the same structure as the encoder but uses different parameters, $\theta^{\mu}_{\mathrm{DC}}$, and recovers the sequence from $\mathbf{v}_{k}$ as
$\tilde{\mathcal{S}}_{k}\triangleq\{\tilde{\mathbf{s}}_{k}(1),\dots,\tilde{\mathbf{s}}_{k}(|\mathcal{A}_{k}|)\}=\mathrm{DCNN}(\mathbf{v}_{k}|\theta_{\mathrm{DC}}),\ \forall k,$ (38)
where $\tilde{\mathcal{S}}_{k}$ is the recovered sequence. In the decoder, each LSTM block in the first layer takes $\mathbf{v}_{k}$ as input, and the recovered sequence is obtained as the collection of output gate vectors from the LSTM blocks in the last layer. The loss function of the SENN is designed to minimize the difference between the original and recovered state sequences using the mean squared error as
$L(\theta^{\mu}_{\mathrm{SE}},\theta_{\mathrm{DC}})=\frac{1}{K}\sum_{k=1}^{K}\sum_{a=1}^{|\mathcal{A}_{k}|}\|\mathbf{s}_{k}(a)-\tilde{\mathbf{s}}_{k}(a)\|_{2}^{2}.$ (39)
The SENN (the encoder NN) and the decoder NN are updated over the above loss functions using stochastic gradient descent with a learning rate $\eta$ for $M$ iterations. Note that only the SENN (the encoder) is used in the NGM NN to embed the state sequences, as shown in Fig. 5, while the decoder is not a component of the NGM NN.
The loss functions of the predictors are designed to minimize the difference between the true values and the predicted ones using binary cross-entropy as
$L(\theta^{\mu}_{\mathrm{PC}})$ $=\sum_{i\neq j}O^{\mathrm{C}}_{i,j}\log_{2}\tilde{O}^{\mathrm{C}}_{i,j}+(1-O^{\mathrm{C}}_{i,j})\log_{2}(1-\tilde{O}^{\mathrm{C}}_{i,j});$ (40)
$L(\theta^{\mu}_{\mathrm{PH}})$ $=\sum_{i\neq j}O^{\mathrm{H}}_{i,j}\log_{2}\tilde{O}^{\mathrm{H}}_{i,j}+(1-O^{\mathrm{H}}_{i,j})\log_{2}(1-\tilde{O}^{\mathrm{H}}_{i,j}).$
The predictors are updated using stochastic gradient descent over the above loss functions with a learning rate $\eta$ for $M$ iterations. Note that both predictor parameters, $\theta^{\mu}_{\mathrm{PC}}$ and $\theta^{\mu}_{\mathrm{PH}}$, are components of the NGM NN, as illustrated in Fig. 5.

### v2

@@ VI-E1 Comparison with PG and DPG
We compare the performance of the ES with the PG and the DPG when a fixed small number of STAs is selected using the batching mode of the DHF, i.e., $K^{\prime}=20$. The performance is measured in Fig. 11 based on the QoS violations, the ratio of approximated optimal slots in (22) to the number of used slots, and the averaged reward performance indicator. The results show that the proposed ES algorithm achieves the lowest QoS violation probability, the least number of slots, and the highest reward performance (approximately $4\sim 10$ times higher), while the PG and DPG algorithms struggle to converge to satisfactory performance. This is because the PG and DPG algorithms require STA-pairwise (or edge-wise) gradients, but no STA-pairwise feedback is available. In comparison, the ES algorithm directly applies the reward feedback in the parameter space to improve the IGL NN.
Figure 11: The performance of ES, PG and DPG for the IGL when $K^{\prime}=20$.
@@ VI-E2 Performance of ES with Different Batching Schemes
We then show the performance of the ES algorithm with and without the batching mode of the DHF. Specifically, we consider 1) the proposed batching method that adaptively increases the number of STAs, $K^{\prime}$, based on the reward performance (with legend “Adp-B”), as shown in Algorithm 2; 2) a batching method that linearly increases $K^{\prime}$ by one in each step (with the legend “Lin-B”); and 3) a method that uses no batching but trains the IGL NN with the total number of STAs, $K=K^{\prime}$ (with the legend “No-B”). Fig. 12 measures the number of STAs, the training time, and the averaged reward performance indicator during the training steps. The results show that the proposed adaptive batching converges to the target number of STAs at approximately $1100$ steps. Additionally, the performance of all three compared methods is similar around $1100$ steps. However, the training time of the proposed adaptive batching method is approximately $4$ times and $2$ times shorter than that of the linear batching method and the no-batching method, respectively. In other words, the proposed method significantly reduces the training time for the IGL. In the remaining simulations, the trained IGL parameters are fixed as those from the last step of the proposed ES with the adaptive batching, without further updates.
VI-F Online Graph Construction with DHF Bucketing Mode
We evaluate the trained IGL NN using the online graph construction process designed in Section V-C with the bucketing mode of DHF when there are $K=1000$ STAs.
Figure 13: The IGL NN in the architecture when STAs are static.
Figure 14: Comparison between the IGL and human intuition-based graphs.
Figure 15: Performance when using different numbers $N_{\text{CH}}$ of channels.
@@ VI-F1 Performance When STAs are Static
Fig. 13 shows the IGL NN’s performance in the cumulative distribution function (CDF) of the number of used slots and the number of QoS violating STAs when the IGL NN operates for different numbers of rounds, $m$, in the architecture. In this architecture, all STA pairs connected by edges in previous rounds are combined into the processing STA pairs for the current round, as shown in (39). The results show that across different rounds, the number of slots to color the generated graph remains consistent, around $30$ slots. It also shows that the IGL NN in the first round results in $2\%$ of STAs experiencing QoS violations. However, by running the online process for more rounds, the QoS violations are fewer than $1\%$. This improvement is due to the bucketing mode in each round, which can miss some contending and interfering STA pairs, as shown in previous simulations in Section VI-D. By adding the previously connected pairs in the IGL process as (39), the likelihood of missing the processing of contending or interfering STA pairs is efficiently reduced.
Fig. 14 compares the number of slots and QoS violations achieved by the IGL with the DHF’s bucketing mode when the number of rounds $m$ is 9 (with legend “IGL-B”), the IGL with all STA pairs processed (with legend “IGL-A”), and the human intuition-based graph models (with legends “CHG” and “IFG”). The results show that the IGL achieves $25\%$ fewer slots than the human intuition-based graph models. This is because the IGL allows slot sharing among STA pairs that are not heavily contending or interfering with each other, while the human intuition-based graph models separate all contending or interfering pairs. Consequently, the human intuition-based graph models achieve fewer QoS violations but use more slots or cause higher delays in periodicity. The results also show that the scheme with partly processed STA pairs using the DHF’s bucketing mode achieves performance similar to the one where the IGL NN processes all STA pairs. This implies that only a subset of STA pairs, e.g., those likely to be contending or interfering, require IGL NN processing, which further suggests that the DHF’s bucketing mode reduces the complexity when constructing the graph.
In practice, Wi-Fi networks often operate on multiple channels to reduce contention and interference. Fig. 15 shows the performance of the IGL NN with the DHF’s bucketing mode when different numbers of channels, $N_{\text{CH}}$, are used. Here, we assume that each neighboring AP operates on a different channel when $N_{\text{CH}}>1$, and each channel has an equal bandwidth of $B=20$ MHz. The results show that using more channels reduces both the number of slots and the number of QoS violations. This is because more channels reduce contention and interference among STAs, allowing more STAs to share the same slot without causing QoS violations. As shown in the results, the performance gain from adding channels eventually plateaus. Increasing the number of channels reduces inter-AP interference because neighboring APs can operate on different channels. Once channel separation among neighboring APs is sufficient, the remaining dominant interference comes mainly from STAs associated with the same AP. Therefore, adding more channels beyond this point provides little additional benefit. We also compare the performance of the IGL with human intuition-based graph models under different numbers of channels. Here, we show only the CHG results, because IFG performs similarly to CHG, as shown previously. The results show that the IGL uses approximately $25\%$ fewer slots than CHG across different numbers of channels, while CHG yields fewer QoS violations than IGL. These results are consistent with the observations for $N_{\text{CH}}=1$.
@@ VI-F2 Performance When STAs are Mobile
We then show the IGL’s performance when STAs are mobile. Specifically, STAs move in a random direction within the simulated area, with speeds varying from 0 to $5$ meters per second, approximately the typical walking speeds of mobile robots. Note that when STAs reach the boundary of the simulated area, they move in another random direction toward the inside of the area. We compare the following schemes: 1) the IGL using the DHF’s bucketing mode and combining previous rounds’ edges (the maximum rounds to combine is $I=20$) into the processed pairs (with legend “IGL-B”); 2) the IGL NN using the DHF’s bucketing mode without combining previous edges in the processed pairs (with legend “No-Comb”); and 3) the IGL NN processing all edge pairs among STAs (with legend “IGL-A”). Fig. 16 measures the average number of slots and the average packet loss rates in the above schemes. The results show that the average number of slots is similar across all schemes (around $30$ slots), while the IGL-B and IGL-A schemes achieve lower packet loss rates. This is because the IGL-B and IGL-A schemes process more contending or hidden STA pairs than the No-Comb scheme. On the other hand, the IGL-B scheme has $30\%$ fewer packet losses than the IGL-A scheme when STAs have high mobility. This is because the DHF’s bucketing mode selectively processes contending or interfering STA pairs, making the IGL process with the DHF’s bucketing mode faster when constructing the graph. Consequently, it responds more timely to the varying locations of STAs. The detailed computing time of the IGL process is measured next. We also compare the performance of the IGL with human intuition-based graph models, CHG and IFG, when STAs are mobile in Fig. 16. The results show that the IGL-B scheme achieves $25\%$ fewer slots than CHG and IFG, while CHG and IFG achieve lower packet loss rates. This is because the human intuition-based graph models separate all contending or interfering STA pairs, while the IGL allows slot sharing among less contending or interfering STA pairs.
@@ VI-G Computational Complexity in Time
We discuss the computational complexity of the proposed methods in terms of the number of STAs, $K$. Training the state embedding (trained over STA-wise), the predictors (trained over STA-pairwise), and the DHF (also trained over STA-pairwise) incurs $\mathcal{O}(K)$, $\mathcal{O}(K^{2})$, and $\mathcal{O}(K^{2})$ complexity, respectively. The batching mode of the DHF requires a $\mathcal{O}(K)$ complexity to compare the random query and collect the batch. Training the EGNN in Algorithm 2 is also processed over STA-pairwise, resulting in an $\mathcal{O}(K^{2})$ complexity. The IGL complexity when constructing/processing the graph depends on the number of STA pairs collected using the bucketing mode of the DHF. Since it is difficult to analyze hash codes/tables generated by DHF, we instead measure the computing time of the online graph construction process quantitatively. Specifically, Fig. 17 compares the computing time of the online process when $K=1000$ in 1) the scheme where the IGL NN only processes STA pairs collected by the DHF bucketing mode (with legend “IGL-B”), and 2) the scheme where the IGL NN processes all STA pairs (with legend “IGL-A”). The results show that the DHF’s bucketing mode reduces the computing time for STA-pairwise operations, such as predictors and edge generation, by a factor of $8$, and the total computing time by a factor of $3$. This is because only the STA pairs in the buckets are processed, while the bucketing mode adds some additional computing time when collecting the STA pairs in the buckets/hash tables. Note that the current hash table construction is single-threaded on a central processing unit, while further time reduction can be achieved by an efficient parallelization on a graphics processing unit.
Figure 16: The IGL NN performance in online architecture with mobile STAs.
Figure 17: The computing time, $\operatorname{\mathbb{E}}[\tau^{(m)}]$, for the online graph constuction process to model and color the graph, including the time spent by the state embedding (Emb.), the DHF’s hash code generation (Hsh.), the DHF’s bucketing mode (Buc.), the predictors (Pre.), edge/graph generation (EG) and the graph coloring (Col.)
@@ VII Conclusion
This paper presented the IGL for managing contention and interference in Wi-Fi 7 networks through coloring-based slot assignment. The proposed IGL framework achieved scalability in terms of network size in training and inference. Specifically, by employing the ES, the IGL eliminated the need for explicit STA-pairwise feedback, whose dimensionality grows quadratically with the number of STAs, enabling stable optimization even as the network scales. In addition, the DHF restricted training and inference to a small subset of likely contending or hidden STA pairs, significantly reducing the computational complexity in large-scale networks. Extensive evaluation demonstrated the scalability and effectiveness of our methods in large Wi-Fi network scenarios. For example, the proposed IGL framework with the DHF reduced the computing time of the online graph construction process by up to a factor of $3$, while achieving $25\%$ fewer slots than human intuition-based graph models with a slight increase in QoS violations.
Future work could focus on further optimizing the graph coloring algorithm and accelerating the computation routines in the IGL. Moreover, the current IGL framework is designed as a centralized approach, where the AP collects the network states from all STAs and generates the interference graph for slot assignment. However, the proposed IGL framework can be extended to a distributed manner by enabling each STA to locally estimate its interference relationships with neighboring STAs based on local observations and limited information exchange. This can further reduce the communication overhead and improve scalability in large-scale networks. Additionally, research could be conducted to explore richer graph representations that integrate the application layer’s context, semantic, and task-oriented information, and to extend the framework to accommodate more heterogeneous network deployments, e.g., with devices with different traffic profiles (see Appendix D for detailed discussions). Furthermore, the proposed IGL framework can be extended to the resource allocation schemes, including OFDMA and multi-user multiple-input multiple-output (MU-MIMO), in Wi-Fi 7 and beyond networks, by designing appropriate graph structures to represent the contention and interference in the schemes. Finally, more efficient implementations of the graph coloring algorithms can be investigated to further balance the load among different slots and reduce the QoS violations, while exploiting the sparsity of the constructed graphs [52].